Supposed the disposable income ranges from 1,110,000 to 1,290,000 and it increases by 30,000 per year. Mpc = .80

When disposable income is 1,000,000 what is the level of consumption is equal to?


1
Expert's answer
2022-02-18T08:40:39-0500

a)MPC=80 ,but "MPC=\frac{Change in consumption}{change in disposable income}"

so "Change in consumption= MPC*Change in disposable income"

"=80*[(1000000)-(1290000-1110000)\n+30000]"

"=63200000 dollars"

but consumption"=a+bY"

Where 'a' is autonomous consumption

b is the MPC

Y is the disposable income

Consumption"=30000+80*1000000"

"=80030000 dollars"

level of consumption"=80030000-63200000"

="16830000 dollars"